OpenStudy (anonymous):

Number (x) 2 -2 3 -3 Square of the number (y) 4 4 9 9 Part A: Does the table represent y as a function of x? Justify your answer. (5 points) Part B: The total cost f(x), in dollars, for renting a rowboat for x hours is shown below: f(x) = 20 + 3x What is the value of f(150), and what does f(150) represent? (5 points) @iheartfood ?

OpenStudy (yttrium):

for part B: f(x) = 20 + 3x f(150) = 20 + 3(150) = 20 + 450 = 470 (Answer) It represents the total cost, in dollars, for renting a rowboat in 150 hours.

OpenStudy (anonymous):

Thank you! did you get part A?

OpenStudy (yttrium):

For part A: No. Proof:  It didn't pass the Vertical Line Test.